find 3 consecutive even integers such that three times the first equals the sum of the other two
Answer by stanbon(75887)

find 3 consecutive even integers such that three times the first equals the sum of the other two
---
1st: 2x-2
2nd: 2x
3rd: 2x+2
---
Equation:
3(2x-2) = 4x+2
6x-12 = 4x+2
2x = 14
x = 7
--------
1st: 2x-2 = 14-2 = 12
2nd: 2x = 14
3rd: 2x+2 = 16
===================
Cheers,
Stan H.
===================